OpenStudy (anonymous):

log (x+4) + log(x+3)= 1

OpenStudy (anonymous):

x^2+7x+12=10 ==>x^2+7x+2=0 ( if the base is 10)

OpenStudy (anonymous):

b^2-4ac=49-8 =41 ==> x=[-7 + ( or -) + sqrt41]/2